Blossom Company bottles and distributes B-Lite, a diet soft drink. The beverage is sold for 50 cents per 16-ounce bottle to retailers, who charge customers 75 cents per bottle. For the year 2020, management estimates the following revenues and costs. 1,500,000 Selling expenses-variable 90,000 Sales Direct materials Direct labor Manufacturing overhead-variable Manufacturing overhead-fixed 410,000 Selling expenses-fixed 330,000 Administrative expenses-variable 20,000 350,000 Administrative expenses-fixed 150,000 50,000 40,000BLOSSOM COMPANY CVP Income Statement (Estimated) For the Year Ending December 31,2020Calculate variable cost per bottle. (Round variable cost per bottle to 3 decimal places, e.g.0.251.) Variable cost per bottle $ eTextbook and Media Compute the break-even point in (1) units and (2) dollars. (Round answers to O decimal places, eg. 1,225.) (1) Compute the break-even point (2) Compute the break-even point $ units eTextbook and Media Compute the contribution margin ratio and the margin of safety ratio. (Round variable cost per bottle to 3 decimal places, eg. 0.25 and final answers to O decimal places, eg, 25%) Contribution margin ratio Margin of safety ratioDetermine the sales dollars required to earn net income of $ 150,000. (Round answer to O decimal places, eg. 1,225.) Required sales dollars $ eTextbook and Media
Answer #1
Blossom Company
CVP Income Statement(Estimated)
Sales 1500000
Variable expenses:
Cost of goods sold 1090000 =410000+330000+350000
Selling expenses 90000
Administrative expenses 20000
Total Variable expenses 1200000
Contribution margin 300000
Fixed expenses:
Cost of goods sold 150000
Selling expenses 50000
Administrative expenses 40000
Total Fixed expenses 240000
Net income 60000
2
Variable cost per bottle 0.40 =1200000/3000000
3
(1) Break-even point 2400000 units =240000/(0.5-0.4)
(2) Break-even point 1200000 =2400000*0.5
4
Contribution margin ratio 20% =300000/1500000
Margin of safety ratio 20% =(1500000-1200000)/1500000
5
Required sales dollars 1950000 =(240000+150000)/20%